OpenStudy (anonymous):

Help Please!!! Direct Variation.... In a circle, the circumference and diameter vary directly. Without using a value of , find the diameter of a circle with a circumference of 154 if you know that in a second circle the diameter is 14 when the circumference is 44.

OpenStudy (anonymous):

@.Sam. can you please help me??? :)

sam (.sam.):

They said that Circumference=diameter By comparing both the circles, rearranging the equation \[\frac{Circumference_1}{Diameter_1}=\frac{Circumference_2}{Diameter_2}\] \[\frac{154}{Diameter_1}=\frac{44}{14}\] \[Diameter_1=49\]